Integrals Giving Inverse Trigonometric Functions


1. ∫(xdx)/√(3-2x-x²)


2. ∫(dy)/(16+y²) from -4 to 4


1
Expert's answer
2021-04-14T13:01:35-0400

1)

"\int{\frac{x}{\sqrt{3-2x-x^2}}dx}"

focus on the denominator

"3-2x-x^2"

"\left(-1\right)\left(x^2+2x-3\right)"

"\left(-1\right)\left(x^2+2x+1-1-3\right)"

"\left(-1\right)\left(\left(x+1\right)^2-4\right)"

"4-\left(x+1\right)^2"

"\int{\frac{x}{\sqrt{4-\left(x+1\right)^2}}dx}"

"\int{\frac{x+1-1}{\sqrt{4-\left(x+1\right)^2}}dx}"

"\int{\frac{x+1}{\sqrt{4-\left(x+1\right)^2}}dx}-\ \int{\frac{1}{\sqrt{4-\left(x+1\right)^2}}dx\ }"

"let\ u=\sqrt{4-\left(x+1\right)^2}"

"du=\ -\frac{2\left(x+1\right)}{\sqrt{4-\left(x+1\right)^2}}dx=>dx=\ -\frac{\sqrt{4-\left(x+1\right)^2}}{2\left(x+1\right)}du"

"\int{\frac{x+1}{\sqrt{4-\left(x+1\right)^2}}dx}=\ -\frac{1}{2}\int d u=-\frac{u}{2}+c=\ -\frac{\sqrt{4-\left(x+1\right)^2}}{2}+c"


"let\ v=x+1=>dv=dx"

"\int{\frac{1}{\sqrt{4-\left(x+1\right)^2}}dx\ }=\ \int{\frac{1}{\sqrt{2^2-v^2}}dv=\sin^{-1}\left(\frac{v}{2}\right)}+d=\sin^{-1}{\left(\frac{x+1}{2}\right)}+d"



"-\frac{\sqrt{4-\left(x+1\right)^2}}{2}-\ \sin^{-1}{\left(\frac{x+1}{2}\right)}+K"


2)

"\int_{-4}^{4}\frac{dy}{16+y^2}"

"\int_{-4}^{4}\frac{dy}{4^2+y^2}"

"\frac{1}{4}\tan^{-1}{\left(\frac{y}{4}\right)}\ \ \left(-4,\ 4\right)"

"\frac{1}{4}\left(\tan^{-1}{\left(1\right)-\tan^{-1}{\left(-1\right)}}\right)"

"\frac{1}{4}\left(\frac{\pi}{4}--\frac{\pi}{4}\right)=\frac{\pi}{8}"